# Limit Laws for Poincaré Recurrence and the Shrinking Target Problem

### v1

@@ 2.4. Almost Sure Invariance Principles for Axiom A Diffeomorphisms
We compare Theorem 2.5 to other known results. The improvement over [13, Corollary 6.2] is two-fold. First, we allow the sequence $(\|\varphi_{n}\|_{\alpha})_{n}$ of Hölder norms to grow at a polynomial rate, and second, we relax the condition on $\delta$ from $\delta>\frac{\sqrt{17}-1}{4}\approx 0.781$ to $\delta>1/2$.
@@ 3.2. Variance estimates for explicit radii
In addition to the above assumptions, suppose that $\mathop{}\!d\mu=h\mathop{}\!dm$ with density $h\in L^{2}(m)$ such that $h\geq c>0$ for some constant $c>0$. Additionally, suppose that the following short returns estimate holds: there exist constants $\delta_{0}>0$ and $p_{0}>1$ such that
@@
for all sufficiently large $k$ and all $l=1,\ldots,(\log k)^{p_{0}}$. If $(r_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,d)$ and $\upsilon>2(\delta_{0}\dim X)^{-1}$, then
@@ 4. Supporting Results
The method for establishing (ASIP) for the shrinking target problem has its roots in the following result by Cuny and Merlevède [10, Theorem 2.3] that establishes (ASIP) for a sequence of random variables $(X_{n})$. We incorporate [10, Remark 2.4] into the statement, which weakens the conditions on boundedness of the sequence $(\mathbf{E}X_{n}^{2})_{n}$. This is the method the authors in [13] employ to establish (ASIP) for the shrinking target problem in the case of non-invertible systems satisfying Condition D.1 with $\|\cdot\|_{\alpha}$ replaced with $\|\cdot\|_{\textup{BV}}$ [13, Theorem 5.1].
@@
Let $(X_{n})$ be a sequence of square-integrable random variables adapted to a non-increasing filtration $(\mathcal{G}_{n})_{n}$. Assume that $\mathbf{E}[X_{n}|\mathcal{G}_{n+1}]=0$ a.s., that $\sigma_{n}^{2}:=\sum_{k=1}^{n}\mathbf{E}[X_{k}^{2}]\to\infty$ and that there exists $s\in[0,1)$ such that $\mathbf{E}[X_{n}^{2}]=\mathcal{O}(\sigma_{n}^{2s})$ for all $n$. Let $(a_{n})_{n}$ be a non-decreasing sequence of positive numbers such that $(a_{n}/\sigma_{n}^{2})_{n}$ is non-increasing and $(a_{n}/\sigma_{n})_{n}$ is non-decreasing. Assume that
@@
The following result is a standard tool for establishing strong Borel–Cantelli lemmas – a non-autonomous version of a strong law of large numbers – by reducing the problem to a second moment analysis. Here it will be used to establish Condition A in Theorem 4.1. We state the result as formulated by Harman [12].
@@
For the shrinking target problem, the hitting sets are of the form $A_{n}=\{x\in X:T^{n}x\in B(y_{n},r_{n})\}$. Since $T$ preserves $\mu$, one obtains that $\mu(A_{n})=\mu(B(y_{n},r_{n}))$. In the recurrence setting, the targets depend on the initial point and one cannot directly use the measure-preserving property. Nevertheless, if the system is rapidly mixing, then one can obtain estimates for the difference between $\mu(E_{n})$ and $\mu(B(x,r_{n}(x)))$, and $\mu(\hat{E}_{n})$ and $\hat{M}_{n}=\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)$. The following lemma was established by Kirsebom, Kunde and Persson [23, Lemma 4.1] for interval systems. In [35, Proposition 4.1], the author establishes (4.1) for uniformly hyperbolic systems. A simple rewriting of the proof of [35, Proposition 4.1], gives (4.2) as well. Note that Kirsebom, Kunde and Persson’s result as it appears in [23] is slightly stronger when restricting to their setting.
@@
In order to leverage Conditions D.2 or D.3, which apply to Hölder continuous observables, we would like to obtain a partition of $X$ for which the indicator function of each set can be well approximated by Hölder continuous observables. This turns out to be difficult in general; however, by removing a subset of $X$ with small measure, Conze and Le Borgne obtain a partition of the remaining space with the desired properties. The statement and proof was originally communicated in private correspondence, but can be found in [34, Lemma 4].
@@
We state a lemma regarding a Hölder continuous approximation of indicator functions of balls as it will be used in many of the proofs. The proof is similar to that of the approximation argument in [23, Lemma 3.1]. Recall the constants $\alpha_{0}$ and $\rho_{0}$ appearing in Condition M.2.
@@ 5. Proof of CLT for Recurrence
We devote this section to the proof of Theorems 2.1 and 2.2. A central argument in the proof of Theorem 2.1 is Lévy’s continuity theorem: the sequence of observables $(f_{n})$ converges in distribution to $f$ if and only if the characteristic functions $\psi_{f_{n}}(t)$ converge to $\psi_{f}(t)$ as $n\to\infty$ for all $t\in\mathbb{R}$ (see e.g. [22, Theorem 5.3]). In particular, we use the fact that if $(f_{n})$ and $(g_{n})$ are sequences of observables such that $g_{n}\to g$ in distribution and $|\psi_{f_{n}}(t)-\psi_{g_{n}}(t)|\to 0$ for all $t\in\mathbb{R}$, then $f_{n}\to g$ in distribution. Another argument that we repeatedly use is the following. Let $f,g$ be two observables. Then, for all $t\in\mathbb{R}$,
@@
By [35, Lemma 3.1, Lemma 3.2], the radii functions $r_{n}\colon X\to(0,\infty)$ are Lipschitz continuous with $\lvert r_{n}\rvert_{\textup{Lip}}=1$ and $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$. Using Condition M.1 together with $M_{n}\gtrsim n^{-\gamma}$, one obtains that $r_{n}(x)\gtrsim n^{-\gamma/s_{0}}$ for all $x\in X$. Furthermore, $\sigma_{n}^{2}\lesssim n^{2}$. Therefore, we may apply Lemma 5.1 with $a_{n}=\sigma_{n}$ and to obtain that
@@
since $\int|S_{n,y}|\mathop{}\!d\mu\leq s_{n}(y)$ by the Cauchy–Schwarz inequality. The assumption that $s_{n}^{2}/\sigma_{n}^{2}\to 1$ in distribution implies that $s_{n}^{2}/\sigma_{n}^{2}\to 1$ in measure. Together with Lemma 3.3, namely that $\|s_{n}^{2}/\sigma_{n}^{2}\|_{L^{1}}\to 1$, one obtains that $s_{n}^{2}/\sigma_{n}^{2}\to 1$ in $L^{1}$ (see e.g. [22, Proposition 4.12]). Since $|s_{n}/\sigma_{n}+1|\geq 1$, we therefore have that
@@
$x\mapsto\frac{1}{\hat{s}_{n}(y)}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(y,r_{k})}(T^{k}x)-\mu(B(y,r_{k}))\bigr),$
where $\hat{\sigma}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{E_{k}})$ and $\hat{s}_{n}^{2}(y)=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k})}\circ T^{k})$. As before, we also assume, without loss of generality, that $r_{n}\leq\rho_{0}$ for all $n\in\mathbb{N}$.
@@
Since $\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)=M_{n}\gtrsim n^{-\gamma}$, we can use Condition M.1 to obtain that $r_{n}\gtrsim n^{-\gamma/s_{0}}$. Clearly, $r_{n}\to 0$ and $\lvert r_{n}\rvert_{\textup{Lip}}=0$. Since $\hat{\sigma}_{n}^{2}\lesssim n^{2}$, we may therefore apply Lemma 5.1 with $a_{n}=\hat{\sigma}_{n}$ to obtain that
@@
$\begin{split}\int_{X}\bigl|\psi_{n}^{\sigma}(t,y)-\psi_{n}^{s}(t,y)\bigr|\mathop{}\!d\mu(y)&\lesssim_{t}\int_{X}\Bigl|\frac{1}{\hat{\sigma}_{n}}-\frac{g}{\hat{s}_{n}(y)}\Bigr|\int_{X}|S_{n,y}(x)|\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\\ &\leq\int_{X}\Bigl|\frac{1}{\hat{\sigma}_{n}}-\frac{g}{\hat{s}_{n}(y)}\Bigr||\hat{s}_{n}(y)|\mathop{}\!d\mu(y)\\ &=\int_{X}\Bigl|\frac{\hat{s}_{n}(y)}{\hat{\sigma}_{n}}-g\Bigr|\mathop{}\!d\mu(y),\end{split}$
@@
The conditions on $h$ imply that $\mu$ satisfies Conditions M.1 and M.2. Furthermore, the assumptions of Lemma 3.6 are satisfied. Hence, $\hat{s}_{n}/\hat{\sigma}_{n}\to g$ in $L^{1}(\mu)$. The assumptions of Theorem 2.2 are therefore satisfied and we conclude. ∎
@@ 6. Proof of ASIP for Non-invertible Systems
We now verify Condition B for $\nu=2$ similarly to how the authors in [13, Theorem 5.2] do. Recall that $a_{k}=\hat{\sigma}_{n}^{2-\varepsilon}$ for some $\varepsilon\in(0,1)$, where $\hat{\sigma}_{n}^{2}=\sum_{k=1}^{n}\mathbf{E}[X_{k}^{2}]$. By (6.1), we have that $\|\tilde{h}_{n}\|_{\infty}\lesssim\log n$. Hence,
@@
To establish Condition B for $\nu=2$, we use the same approach as employed in [13, Theorem 5.1]. Notice that, just as in the proof of Theorem 2.3, we have that $\hat{\sigma}_{n}^{2}\gtrsim\sigma_{n}^{2}\gtrsim n^{\delta}$. Thus, since $a_{n}^{2}=\hat{\sigma}_{n}^{4-2\varepsilon}\gtrsim n^{(2-\varepsilon)\delta}$,
@@ 7. Proof of ASIP for Axiom A Diffeomorphisms
Recall that in the statement of Theorem 2.5, $T\colon X\to X$ is an Axiom A diffeomorphism restricted to a mixing component and $\mu$ is a Gibbs measure corresponding to a Hölder continuous observable on $X$. In this setting, it is well known (see [4]) that there exists a (finite) Markov partition of $X$ that allows one to encode $(X,T,\mu)$ as a sub-shift of finite type (SFT), which we denote as $(\Sigma,\sigma,\mu^{*})$. Let $\pi\colon\Sigma\to X$ denote the quotient map corresponding to the aforementioned choice of Markov partition. Then $\pi\circ\sigma=T\circ\pi$ and $\mu=\pi^{*}\mu^{*}$. Recall as well that $(\varphi_{n})$ is a sequence of $\alpha$-Hölder continuous functions such that $\|\varphi_{n}\|_{\infty}\lesssim(\log n)^{q}$ and $\|\varphi_{n}\|_{\alpha}\lesssim n^{p}$ for some constants $p,q>0$. We note the following result as stated by Bowen [4, Lemma 4.1], but it is due to Hirsch and Pugh [16, p. 140].
@@
Our first lemma is a modification of the ‘Sinai trick’ (explained in [4, 29] for instance), which allows one to reduce proofs of certain properties of two-sided SFTs (such as existence of Gibbs measures) to those of one-sided SFTs. The usual Sinai trick is employed to obtain, from a function $f\colon\Sigma\to\mathbb{R}$, a function $f^{+}\colon\Sigma\to\mathbb{R}$ depending only on future coordinates, i.e., $f^{+}(\underline{x})=f^{+}(\underline{y})$ whenever $\underline{x},\underline{y}\in\Sigma$ with $x_{i}=y_{i}$ for all $i\geq 0$. The modification we make accounts for the non-stationary aspect of our problem; we consider a sequence of functions $(\varphi_{n})$ as opposed just a single function $\varphi$. In the case that $(\varphi_{n})$ is uniformly bounded in $\|\cdot\|_{\alpha}$, Haydn et al [13, Corollary 6.2] prove a modified Sinai trick. In our case, in order to apply our results to the shrinking target problem, we do not assume that the sequence $(\varphi_{n})$ is uniformly bounded in $\|\cdot\|_{\alpha}$. We also keep track of the expectation and variance corresponding to the new sequence of functions. Once we establish our modified Sinai trick, we can reduce the problem to the one-sided sub-shift and then use Conditions D.1, M.1 and M.2 since they hold for $(\Sigma^{+},\sigma,\mu^{*})$.
@@
We modify the proof of the usual Sinai trick (see for instance [4, 29]) and incorporate a modification due to [13, Corollary 6.2].

### v2

@@ 2.4. Almost Sure Invariance Principles for Axiom A Diffeomorphisms
We compare Theorem 2.5 to other known results. The improvement over [14, Corollary 6.2] is two-fold. First, we allow the sequence $(\|\varphi_{n}\|_{\alpha})_{n}$ of Hölder norms to grow at a polynomial rate, and second, we relax the condition on $\delta$ from $\delta>\frac{\sqrt{17}-1}{4}\approx 0.781$ to $\delta>1/2$.
@@ 3.2. Variance estimates for explicit radii
In addition to the above assumptions, suppose that $\mathop{}\!d\mu=h\mathop{}\!dm$ with density $h\in L^{q}(m)$ for some $q>2$ such that $h\geq c>0$ for some constant $c>0$. Additionally, suppose that the following short returns estimate holds: there exist constants $\delta_{0}>0$ and $p_{0}>1$ such that
@@
for all sufficiently large $k$ and all $l=1,\ldots,(\log k)^{p_{0}}$. If $(r_{n})$ satisfies Condition $\mathcal{S}_{\gamma,\upsilon}$ for some $\gamma\in(0,\dim X)$ and $\upsilon>2(\delta_{0}\dim X)^{-1}$, then
@@ 4. Supporting Results
The method for establishing (ASIP) for the shrinking target problem has its roots in the following result by Cuny and Merlevède [10, Theorem 2.3] that establishes (ASIP) for a sequence of random variables $(X_{n})$. We incorporate [10, Remark 2.4] into the statement, which weakens the conditions on boundedness of the sequence $(\mathbf{E}X_{n}^{2})_{n}$. This is the method the authors in [14] employ to establish (ASIP) for the shrinking target problem in the case of non-invertible systems satisfying Condition D.1 with $\|\cdot\|_{\alpha}$ replaced with $\|\cdot\|_{\textup{BV}}$ [14, Theorem 5.1].
@@
Let $(X_{n})$ be a sequence of square-integrable random variables adapted to a non-increasing filtration $(\mathcal{G}_{n})_{n}$ (i.e., $X_{n}$ is $\mathcal{G}_{n}$-measurable for all $n\in\mathbb{N}$). Assume that $\mathbf{E}[X_{n}|\mathcal{G}_{n+1}]=0$ a.s., that $\sigma_{n}^{2}:=\sum_{k=1}^{n}\mathbf{E}[X_{k}^{2}]\to\infty$ and that there exists $s\in[0,1)$ such that $\mathbf{E}[X_{n}^{2}]=\mathcal{O}(\sigma_{n}^{2s})$ for all $n$. Let $(a_{n})_{n}$ be a non-decreasing sequence of positive numbers such that $(a_{n}/\sigma_{n}^{2})_{n}$ is non-increasing and $(a_{n}/\sigma_{n})_{n}$ is non-decreasing. Assume that
@@
The following result is a standard tool for establishing strong Borel–Cantelli lemmas – a non-autonomous version of a strong law of large numbers – by reducing the problem to a second moment analysis. Here it will be used to establish Condition A in Theorem 4.1. We state the result as formulated by Harman [13, Lemma 1.5].
@@
For the shrinking target problem, the hitting sets are of the form $A_{n}=\{x\in X:T^{n}x\in B(y_{n},r_{n})\}$. Since $T$ preserves $\mu$, one obtains that $\mu(A_{n})=\mu(B(y_{n},r_{n}))$. In the recurrence setting, the targets depend on the initial point and one cannot directly use the measure-preserving property. Nevertheless, if the system is rapidly mixing, then one can obtain estimates for the difference between $\mu(E_{n})$ and $\mu(B(x,r_{n}(x)))$, and $\mu(\hat{E}_{n})$ and $\hat{M}_{n}=\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)$. The following lemma was established by Kirsebom, Kunde and Persson [25, Lemma 4.1] for interval systems. In [38, Proposition 4.1], the author establishes (4.1) for uniformly hyperbolic systems. A simple rewriting of the proof of [38, Proposition 4.1] gives (4.2) as well. Note that Kirsebom, Kunde and Persson’s result as it appears in [25] is slightly stronger when restricting to their setting.
@@
In order to leverage Conditions D.2 or D.3, which apply to Hölder continuous observables, we would like to obtain a partition of $X$ for which the indicator function of each set can be well approximated by Hölder continuous observables. This turns out to be difficult in general; however, by removing a subset of $X$ with small measure, Conze and Le Borgne obtain a partition of the remaining space with the desired properties. The statement and proof was originally communicated in private correspondence, but can be found in [36, Lemma 4]. This lemma fixes a slight inaccuracy in the proof of Theorem 3.3 in Conze–Le Borgne [9].
@@
We state a lemma regarding a Hölder continuous approximation of indicator functions of balls as it will be used in many of the proofs. The proof is similar to that of the approximation argument in [25, Lemma 3.1]. Recall the constants $\alpha_{0}$ and $\rho_{0}$ appearing in Condition M.2.
@@ 5. Proof of CLT for Recurrence
We devote this section to the proof of Theorems 2.1 and 2.2. A central argument in the proof of Theorem 2.1 is Lévy’s continuity theorem: the sequence of observables $(f_{n})$ converges in distribution to $f$ if and only if the characteristic functions $\psi_{f_{n}}(t)$ converge to $\psi_{f}(t)$ as $n\to\infty$ for all $t\in\mathbb{R}$ (see e.g. [24, Theorem 5.3]). In particular, we use the fact that if $(f_{n})$ and $(g_{n})$ are sequences of observables such that $g_{n}\to g$ in distribution and $|\psi_{f_{n}}(t)-\psi_{g_{n}}(t)|\to 0$ for all $t\in\mathbb{R}$, then $f_{n}\to g$ in distribution. Another argument that we repeatedly use is the following. Let $f,g$ be two observables. Then, for all $t\in\mathbb{R}$,
@@
By [38, Lemma 3.1, Lemma 3.2], the radii functions $r_{n}\colon X\to(0,\infty)$ are Lipschitz continuous with $\lvert r_{n}\rvert_{\textup{Lip}}=1$ and $r_{n}\to 0$ uniformly on $\operatorname{supp}\mu$. Using Condition M.1 together with $M_{n}\gtrsim n^{-\gamma}$, one obtains that $r_{n}(x)\gtrsim n^{-\gamma/s_{0}}$ for all $x\in X$. Furthermore, $\sigma_{n}^{2}\lesssim n^{2}$. Therefore, we may apply Lemma 5.1 with $a_{n}=\sigma_{n}$ and to obtain that
@@
since $\int|S_{n,y}|\mathop{}\!d\mu\leq s_{n}(y)$ by the Cauchy–Schwarz inequality. The assumption that $s_{n}^{2}/\sigma_{n}^{2}\to 1$ in distribution implies that $s_{n}^{2}/\sigma_{n}^{2}\to 1$ in measure. Together with Lemma 3.3, namely that $\|s_{n}^{2}/\sigma_{n}^{2}\|_{L^{1}}\to 1$, one obtains that $s_{n}^{2}/\sigma_{n}^{2}\to 1$ in $L^{1}$ (see e.g. [24, Proposition 4.12]). Since $|s_{n}/\sigma_{n}+1|\geq 1$, we therefore have that
@@
$x\mapsto\frac{g(y)}{\hat{s}_{n}(y)}\sum_{k=1}^{n}\bigl(\mathbf{1}_{B(y,r_{k})}(T^{k}x)-\mu(B(y,r_{k}))\bigr),$
where $\hat{\sigma}_{n}^{2}=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{\hat{E}_{k}})$ and $\hat{s}_{n}^{2}(y)=\operatorname{Var}(\sum_{k=1}^{n}\mathbf{1}_{B(y,r_{k})}\circ T^{k})$. As before, we also assume, without loss of generality, that $r_{n}\leq\rho_{0}$ for all $n\in\mathbb{N}$.
@@
Since $\int\mu(B(x,r_{n}))\mathop{}\!d\mu(x)=\hat{M}_{n}\gtrsim n^{-\gamma}$, we can use Condition M.1 to obtain that $r_{n}\gtrsim n^{-\gamma/s_{0}}$. Clearly, $r_{n}\to 0$ and $\lvert r_{n}\rvert_{\textup{Lip}}=0$. Since $\hat{\sigma}_{n}^{2}\lesssim n^{2}$, we may therefore apply Lemma 5.1 with $a_{n}=\hat{\sigma}_{n}$ to obtain that
@@
$\begin{split}\int_{X}\bigl|\psi_{n}^{\sigma}(t,y)-\psi_{n}^{s}(t,y)\bigr|\mathop{}\!d\mu(y)&\lesssim_{t}\int_{X}\Bigl|\frac{1}{\hat{\sigma}_{n}}-\frac{g(y)}{\hat{s}_{n}(y)}\Bigr|\int_{X}|S_{n,y}(x)|\mathop{}\!d\mu(x)\mathop{}\!d\mu(y)\\ &\leq\int_{X}\Bigl|\frac{1}{\hat{\sigma}_{n}}-\frac{g(y)}{\hat{s}_{n}(y)}\Bigr||\hat{s}_{n}(y)|\mathop{}\!d\mu(y)\\ &=\int_{X}\Bigl|\frac{\hat{s}_{n}(y)}{\hat{\sigma}_{n}}-g(y)\Bigr|\mathop{}\!d\mu(y),\end{split}$
@@
The conditions on $h$ imply that $\mu$ satisfies Conditions M.1 and M.2. Furthermore, the assumptions of Lemma 3.6 are satisfied. Hence, $\hat{s}_{n}/\hat{\sigma}_{n}\to g$ in $L^{1}(\mu)$ for $g=\sqrt{h}/\sqrt{\mu(h)}$. The assumptions of Theorem 2.2 are therefore satisfied and we conclude. ∎
@@ 6. Proof of ASIP for Non-invertible Systems
We now verify Condition B for $\nu=2$ similarly to how the authors in [14, Theorem 5.2] do. Recall that $a_{k}=\hat{\sigma}_{n}^{2-\varepsilon}$ for some $\varepsilon\in(0,1)$, where $\hat{\sigma}_{n}^{2}=\sum_{k=1}^{n}\mathbf{E}[X_{k}^{2}]$. By (6.1), we have that $\|\tilde{h}_{n}\|_{\infty}\lesssim\log n$. Hence,
@@
To establish Condition B for $\nu=2$, we use the same approach as employed in [14, Theorem 5.1]. Notice that, just as in the proof of Theorem 2.3, we have that $\hat{\sigma}_{n}^{2}\gtrsim\sigma_{n}^{2}\gtrsim n^{\delta}$. Thus, since $a_{n}^{2}=\hat{\sigma}_{n}^{4-2\varepsilon}\gtrsim n^{(2-\varepsilon)\delta}$,
@@ 7. Proof of ASIP for Axiom A Diffeomorphisms
Recall that in the statement of Theorem 2.5, $T\colon X\to X$ is an Axiom A diffeomorphism restricted to a mixing component and $\mu$ is a Gibbs measure corresponding to a Hölder continuous observable on $X$. In this setting, it is well known (see [4]) that there exists a (finite) Markov partition of $X$ that allows one to encode $(X,T,\mu)$ as a sub-shift of finite type (SFT), which we denote as $(\Sigma,\sigma,\mu^{*})$. Let $\pi\colon\Sigma\to X$ denote the quotient map corresponding to the aforementioned choice of Markov partition. Then $\pi\circ\sigma=T\circ\pi$ and $\mu=\pi^{*}\mu^{*}$. Recall as well that $(\varphi_{n})$ is a sequence of $\alpha$-Hölder continuous functions such that $\|\varphi_{n}\|_{\infty}\lesssim(\log n)^{q}$ and $\|\varphi_{n}\|_{\alpha}\lesssim n^{p}$ for some constants $p,q>0$. We note the following result as stated by Bowen [4, Lemma 4.1], but it is due to Hirsch and Pugh [18, p. 140].
@@
Our first lemma is a modification of the ‘Sinai trick’ (explained in [4, 31] for instance), which allows one to reduce proofs of certain properties of two-sided SFTs (such as existence of Gibbs measures) to those of one-sided SFTs. The usual Sinai trick is employed to obtain, from a function $f\colon\Sigma\to\mathbb{R}$, a function $f^{+}\colon\Sigma\to\mathbb{R}$ depending only on future coordinates, i.e., $f^{+}(\underline{x})=f^{+}(\underline{y})$ whenever $\underline{x},\underline{y}\in\Sigma$ with $x_{i}=y_{i}$ for all $i\geq 0$. The modification we make accounts for the non-stationary aspect of our problem; we consider a sequence of functions $(\varphi_{n})$ as opposed just a single function $\varphi$. In the case that $(\varphi_{n})$ is uniformly bounded in $\|\cdot\|_{\alpha}$, Haydn et al [14, Corollary 6.2] prove a modified Sinai trick. In our case, in order to apply our results to the shrinking target problem, we do not assume that the sequence $(\varphi_{n})$ is uniformly bounded in $\|\cdot\|_{\alpha}$. We also keep track of the expectation and variance corresponding to the new sequence of functions. Once we establish our modified Sinai trick, we can reduce the problem to the one-sided sub-shift and then use Condition D.1 since it holds for $(\Sigma^{+},\sigma,\mu^{*})$ (see [4]).
@@
We modify the proof of the usual Sinai trick (see for instance [4, 31]) and incorporate a modification due to [14, Corollary 6.2].